OpenStudy (anonymous):

2|3x – 5| – 8 = 8 Solve the absolute value equation. Graph the solution. on a number line

OpenStudy (anonymous):

Hint: IxI = x for x>=0 and -x for x<0

OpenStudy (anonymous):

does that help?

OpenStudy (anonymous):

@sueprbuster4 ?????/

OpenStudy (anonymous):

abs%282x%2B4%29=8 Start with the given equation Break up the absolute value (remember, if you have abs%28x%29=a, then x=-a or x=a) 2x%2B4=-8 or 2x%2B4=8 Set the expression 2x%2B4 equal to the original value 8 and it's opposite -8 Now lets focus on the first equation 2x%2B4=-8 2x=-8-4Subtract 4 from both sides 2x=-12 Combine like terms on the right side x=%28-12%29%2F%282%29 Divide both sides by 2 to isolate x x=-6 Divide Now lets focus on the second equation 2x%2B4=8 2x=8-4Subtract 4 from both sides 2x=4 Combine like terms on the right side x=%284%29%2F%282%29 Divide both sides by 2 to isolate x x=2 Divide So the solutions to abs%282x%2B4%29=8 are: x=-6 or x=2 To graph these on the number line, just plot points at -6 and 2 on the number line. If you need more help, email me at jim_thompson5910@hotmail.com Also, feel free to check out my website. Jim Answer by stanbon(51684) About Me (Show Source): You can put this solution on YOUR website! absolute value sign 2x+4 absolute value sign =8 -------------- |2x+4| = 8 --- That thing inside the bars must be 8 or -8

OpenStudy (anonymous):

Solve 2x+4=8 or 2x+4 = -8 --- 2x = 4 or 2x = -12 x = 2 or x = -6

OpenStudy (anonymous):

sorry went completely afk let me read everything

OpenStudy (anonymous):

ok

OpenStudy (theviper):

@abayomi12 so nice answer :)

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

alright thanks

OpenStudy (anonymous):

ok